Please help me solve a, b, c, and d

5. (FC Debt): Suppose EAC in Thailand took out a $6 million loan at the interest rate of 8.0% and the spot exchange rate was B41.0/S when the loan was initiated. EAC had to repay the loan plus interest payment in one year (a) If the spot rate in one year is B44.0/S, what is the foreign exchange loss (in Thai baht) incurred on the loan of S6 million due to the change in exchange rate? (b) What is the effective cost of funds in Thai baht terms in part (a)? (c) Assume Thai government wishes to push the $ value of the Thai baht down by 10% from B41.0/S over the coming year to promote its export competitiveness in dollar markets, what is the new exchange rate following the devaluation of 10%? (c)? (d) What is the effective cost of funds in Thai baht terms based on the devaluation of 10% in part (Note: the effective cost of funds-(Thai baht needed to pay in one year)/(the current loan in Thai baht terms)-1)
Answer #1

1.
=6*44*(1+8%)-6*41*(1+8%)=19.44 million or 19440000

2.
=(6*44*(1+8%)/(6*41))-1=15.902439%

3.
=41/0.9=45.56

4.
=(6*41/0.9*(1+8%)/(6*41))-1=20%